Two neutral particles are kept 1 m apart. Suppose by some mechanism some charge is transferred from one particle to the other and the electric potential energy lost is completely converted into a photon. Calculate the longest and the next smaller wavelength of the photon possible

Grade:12

1 Answers

Arun
6 years ago
r = 1 m Energy = kq^2/R = kq^2/1 Now, kQ^2/1 = hc/λ or λ = hc/kq^2 For max ‘λ’, ‘q’ should be min, For λ hc/kq^2 – 0.863 * 10^3 = 863 m. For next smaller wavelength = 6.63 * 3 *10^-34 * 10^8/9 * 10^9 *(1.6 * 2)^2 * 10^-38 = 863/4 = 215.74 m